Differentiate the function.
g(x) = ln(x sqrt(x^2 - 7))
1
Expert's answer
2011-02-28T05:36:06-0500
g'(x) = (x √(x2-7))'/ x√(x2-7) = [√(x2-7) +x*2x/(2√(x2-7))]/ x√(x2-7) = [ x2 - 7 + x2]/x(x2-7) = (2x2-7) / x(x2-7)